# Meta-Pretraining for Zero-Shot Cross-Lingual Named Entity Recognition in Low-Resource Philippine Languages

### v1

@@ Abstract
David Demitri Africathanks: Corresponding author: dda28@cam.ac.uk     Suchir Salhan     Yuval Weiss Paula Buttery     Richard Diehl Martinez University of Cambridge
@@ 1 Introduction
Voice system ✓ Four-way ✓ Reduced two-way
Case marking ✓ Obligatory ✗ Often dropped
Borrowing / code-switch ✓ High density ✗ More conservative
Morphological richness ✓ Productive affixation ✓ Regular affixation
Word order flexibility
Pronominal systems ✓ Rich clitic pronouns ✓ Similar
Reduplication ✓ Common ✓ Widespread
Orthography variation ✓ Multiple conventions ✗ Multiple conventions
Pivot marking ✓ Consistently overt ✓ Overt but less consistent
@@ 2.1 Motivation
Meta-learning addresses this by shaping initializations for quick adaptation. Model-Agnostic Meta-Learning (MAML) optimizes an LM backbone so that a few gradient steps yield high performance on a new task (Finn et al., 2017). We ask whether such an initialization, learned entirely without Tagalog/Cebuano exposure, can transfer to these languages’ distinct morphological and lexical cues for NER. Prior NLP studies have tested this mostly on English or on “BERT-scale” encoder models (Wu et al., 2020; Ma et al., 2022; Li et al., 2020; de Lichy et al., 2021); we explore whether episodic meta-pretraining of small decoder LMs, without any exposure to Tagalog or Cebuano, can still yield zero-shot gains for NER. We do not evaluate a multilingual language-model baseline, as our objective is to isolate the effect of episodic meta-pretraining under a matched corpus and schedule; training a competitive multilingual baseline would require different data and budgets, confounding a like-for-like comparison.
@@ 2.6 Baselines
For each capacity tier we also evaluate a "vanilla" Pico model (no MAML, pure autoregressive loss) under identical data, schedule and compute. Pretraining results can be found in Appendix E with model configuration details in Appendix F.
@@ 4.2 Tag-level Analysis
Figure 6 reports per-entity F1 obtained after head-only finetuning the Slovak CRF head on each MAML checkpoint. PER climbs to 0.6-0.7 while LOC and ORG remain at zero. This is not a case of the classifier “over-fitting” in the usual sense—i.e. collapsing to always predicting a single label. A linear-chain CRF is free to emit any BIO tag at any position; if it were truly degenerate we would see train loss stagnate near the log-uniform baseline and the PER curve itself would also be flat. Instead, train loss converges to the same narrow band for every checkpoint (Fig.4) and PER performance tracks the amount of meta-pretraining, so the head is learning a genuine decision boundary. It simply has informative features for people but none for locations or organisations.
@@ 4.3 Word-level Analysis
Figures 7(a)7(d) visualise the checkpoint-by-checkpoint evolution of token-level confidence (p(correct tag)p($\text${correct tag})) for the ten most frequent surface words in each evaluation set. Entities and non-entities are split so the dynamic range is not drowned out by O tokens. Two qualitative patterns emerge.
@@ 6 Related Work
Although most work applies meta-learning at fine-tuning time, a growing line of research embeds meta-objectives directly into pretraining. Raghu et al. (2021) showed that framing parameter-efficient adapter learning as a bilevel problem yields representations that fine-tune more effectively than standard PEFT. Hou et al. (2022) extend this to full transformers. Miranda et al. (2023) argue that explicit MAML objectives can outperform fixed pretraining on highly diverse task distributions. Ke et al. (2021) integrate a MAML-style inner loop into a multi-criteria Chinese Word Segmentation pretraining task.
@@ 7 Conclusion
All training runs stop at exactly six thousand outer steps, a horizon that may be too short for the largest model, so the conclusions derived only cover a fraction of the training budget a corporate setup might have. A more diverse and multilingual corpus may alter both quantitative and qualitative conclusions, and varying languages in the meta-task is a natural way to extend this work. Qualitative analysis was conducted on a single configuration and single seed due to cost and GPU constraints. Qualitative analysis was conducted by a native Tagalog speaker with a register typical of Manila, and a wide variety of perspectives would improve the robustness of the analysis. Finally (and most naturally), our focus on only two Austronesian languages controls for certain lexical and syntactic divergences but limits the generality of the typological conclusions; extending to a broader set of Philippine and Malayo-Polynesian languages is a natural next step.
@@ Appendix A NER-Relevant Typological Features of Cebuano and Tagalog
Figure 8: Surface cues for named entities. Tagalog typically provides an overt personal article (si/ni) before names; Cebuano may show the same article, but zero-marked variants also occur in some registers/contexts, reducing overt anchors.
@@ Appendix C Pseudocode
2:initialize inner-optimizer SGD on head $h_{\phi}$
@@
10:   // Meta-learning episode
@@
26:  // outer-step and logging
@@
14:  // outer-step when accumulated

### v2

@@ Abstract
David Demitri Africathanks: Corresponding author: david.demitri.africa@gmail.com     Suchir Salhan     Yuval Weiss Paula Buttery     Richard Diehl Martinez University of Cambridge
@@ 1 Introduction
[…]
@@ 2.1 Motivation
Meta-learning addresses this by shaping initializations for quick adaptation. Model-Agnostic Meta-Learning (MAML) optimizes an LM backbone so that a few gradient steps yield high performance on a new task (Finn et al., 2017). We ask whether such an initialization, learned entirely without Tagalog/Cebuano exposure, can transfer to these languages’ distinct morphological and lexical cues for NER. Our working hypothesis is that a pretraining routine that is itself optimized for rapid adaptation will induce representations that generalize more readily across languages. Prior NLP studies have tested this mostly on English or on “BERT-scale” encoder models (Wu et al., 2020; Ma et al., 2022; Li et al., 2020; de Lichy et al., 2021); we explore whether episodic meta-pretraining of small decoder LMs, without any exposure to Tagalog or Cebuano, can still yield zero-shot gains for NER. We do not evaluate a multilingual language-model baseline, as our objective is to isolate the effect of episodic meta-pretraining under a matched corpus and schedule; training a competitive multilingual baseline would require different data and budgets, confounding a like-for-like comparison.
@@ 2.6 Baselines
For each capacity tier we also evaluate a "vanilla" Pico model (no MAML, pure autoregressive loss) under identical data, schedule and compute. Pretraining results can be found in Appendix E with model configuration details in Appendix F. A more detailed discussion of pretraining results and overall methodology can be found in Africa et al. (2025).
@@ 4.2 Tag-level Analysis
Figure 6 reports per-entity F1 obtained after head-only finetuning the Slovak CRF head on each MAML checkpoint. PER climbs to 0.6-0.7 while LOC and ORG remain at zero. This is not a case of the classifier “over-fitting” in the usual sense—i.e. collapsing to always predicting a single label. A linear‐chain CRF is free to emit any BIO tag at any position; if it were truly degenerate we would see train loss stagnate near the log-uniform baseline and the PER curve itself would also be flat. Instead, train loss converges to the same narrow band for every checkpoint (Fig.4) and PER performance tracks the amount of meta-pretraining, so the head is learning a genuine decision boundary. It simply has informative features for people but none for locations or organisations.
@@ 4.3 Word-level Analysis
Figures 7(a)7(d) visualise the checkpoint‐by‐checkpoint evolution of token-level confidence (p(correct tag)p($\text${correct tag})) for the ten most frequent surface words in each evaluation set. Entities and non-entities are split so the dynamic range is not drowned out by O tokens. Two qualitative patterns emerge.
@@ 6 Related Work
Although most work applies meta‐learning at fine‐tuning time, a growing line of research embeds meta‐objectives directly into pretraining. Raghu et al. (2021) showed that framing parameter-efficient adapter learning as a bilevel problem yields representations that fine‐tune more effectively than standard PEFT. Hou et al. (2022) extend this to full transformers. Miranda et al. (2023) argue that explicit MAML objectives can outperform fixed pretraining on highly diverse task distributions. Ke et al. (2021) integrate a MAML‐style inner loop into a multi-criteria Chinese Word Segmentation pretraining task.
@@ 7 Conclusion
The gains are most pronounced for person entities and head-only finetuning, and scale best with larger model capacities. All training runs stop at exactly six thousand outer steps, a horizon that may be too short for the largest model, so the conclusions derived only cover a fraction of the training budget a corporate setup might have. A more diverse and multilingual corpus may alter both quantitative and qualitative conclusions, and varying languages in the meta-task is a natural way to extend this work. Qualitative analysis was conducted on a single configuration and single seed due to cost and GPU constraints. Qualitative analysis was conducted by a native Tagalog speaker with a register typical of Manila, and a wide variety of perspectives would improve the robustness of the analysis. Finally (and most naturally), our focus on only two Austronesian languages controls for certain lexical and syntactic divergences but limits the generality of the typological conclusions; extending to a broader set of Philippine and Malayo-Polynesian languages is a natural next step.
@@ Appendix A NER-Relevant Typological Features of Cebuano and Tagalog
Figure 8: Surface cues for named entities. Tagalog typically provides an overt personal article (si/ni) before names; Cebuano may show the same article, but zero‑marked variants also occur in some registers/contexts, reducing overt anchors.
@@ Appendix C Pseudocode
2:initialize inner‐optimizer SGD on head $h_{\phi}$
@@
10:   // Meta‐learning episode
@@
26:  // outer‐step and logging
@@
14:  // outer‐step when accumulated